# Study of $1^{--}$ P wave charmoniumlike and bottomoniumlike tetraquark spectroscopy

### v2

@@ Abstract
The masses of $1^{--}$ P-wave charmonium-like and bottomonium-like tetraquark states are calculated in a constituent quark model (CQM) where the Cornell-like potential and Breit-Fermi interaction are employed. All model parameters were imported from previous work, and predetermined by studying the low-lying conventional S- and P-wave light, charmed, bottom, charmonium, and bottomonium meson mass spectra. The decay widths of $1^{--}$ P-wave tetraquark states are calculated for possible two-body strong decay channels within the rearrangement mechanism, including $\omega\chi_{cJ}$, $\eta J/\psi$, and $\rho J/\psi$ for charmonium-like tetraquarks, and $\omega\chi_{bJ}$ for bottomonium-like tetraquarks. The tetraquark theoretical results are compared with the selected exotic states, also known as Y states, and tentative assignments are suggested. This study suggests that $\psi(4230)$, $\psi(4360)$, $\psi(4660)$, and $\Upsilon$(10753) may be P-wave tetraquark states and that multiple states might exist around 4.23 GeV and 4.36 GeV.
@@ II.1 Quark configurations of tetraquark
The construction of tetraquark states follows to the principle that a tetraquark state must be a color singlet, meaning the color wave function of the tetraquark must be a $[222]_{1}$ singlet within the $SU_{c}(3)$ group. In this work we consider only charmonium-like and bottomonium-like ($q_{1}Q_{2}\bar{q}_{3}\bar{Q}_{4}$) states, where $q_{1}$ and $Q_{2}$ are light quark and heavy quark, and $\bar{q}_{3}$ and $\bar{Q}_{4}$ are light antiquark and heavy antiquark respectively.
The permutation symmetry of the two-quark cluster ($qQ$) in tetraquark states is described by the Young tabloids $[2]_{6}$ and $[11]_{\bar{3}}$ of the $SU_{c}(3)$ group, while the color configuration of the two-antiquark cluster ($\bar{q}\bar{Q}$) consists of a $[211]_{3}$ triplet and a $[22]_{\bar{6}}$ antisextet.
Thus, a $[222]_{1}$ color singlet for tetraquark states demands the following configurations: $[2]_{6}(q_{1}Q_{2})\otimes[22]_{\bar{6}}(\bar{q}_{3}\bar{Q}_{4})$ and $[11]_{\bar{3}}(q_{1}Q_{2})\otimes[211]_{3}(\bar{q}_{3}\bar{Q}_{4})$, which correspond to the color sextet-antisextet ($6_{c}\otimes\bar{6}_{c}$) and triplet-antitriplet ($\bar{3}_{c}\otimes 3_{c}$) configurations, respectively. The explicit color wave function for each tetraquark color configuration can be found in the previous work Zhao et al. (2021).
@@
×$\times\psi_{n_{\chi_{1}}l_{\chi_{1}}}(\vec{\chi}_{1}\,)\otimes\psi_{n_{\chi_{2}}l_{\chi_{2}}}(\vec{\chi}_{2}\,)\otimes\psi_{n_{\chi_{3}}l_{\chi_{3}}}(\vec{\chi}_{3}),$
@@ II.2 Hamiltonian
$\vec{L_{ij}}=\vec{r_{ij}}\times\vec{p_{ij}}=\vec{L_{ij}}\times\frac{m_{i}\vec{p_{i}}-m_{j}\vec{p_{j}}}{m_{i}+m_{j}}.$ (8)
@@
By importing the coupling constants $a$, $b$, and $B_{0}$, as well as the constituent quark masses $m_{u}$, $m_{c}$, and $m_{b}$ from previous work Zhao et al. (2021), the mass spectra of S- and P-wave light, charmed, bottom, charmonium, and bottomonium conventional mesons are calculated in the Hamiltonian in Eq. (II.2). A comparison of the theoretical results, presented in Table 1, with experimental data from the PDG Navas et al. (2024), yields a model coupling parameter $\sigma_{0}=0.85$.
@@ II.3 Two-body strong decay
The two-body strong decay properties of the bound states, specifically the Y tetraquark states within the rearrangement mechanism, are studied, as shown in Figure 1. The spin-color factors of the transition amplitudes $T_{cs}$ in the two-body strong decay channels take the form:
@@ III.1 Results
The predetermined as well as imported parameters are applied to predict the mass of Y tetraquark states in the Hamiltonian in Eq. (II.2) including the color-spin interaction $V_{ss}$ which may mix up different color-spin configurations.
Because of the cross terms are listed in Appendix B, eigenstates of the Hamiltonian for S=0 are linear combinations of $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(0\otimes 0)}$, $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(1\otimes 1)}$, $\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(0\otimes 0)}$, and $\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(1\otimes 1)}$. Eigenstates for S=2 are linear combinations of $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=2}_{(1\otimes 1)}$ and $\psi^{c}_{6\otimes\bar{6}}\psi^{S=2}_{(1\otimes 1)}$. There is no configuration mixing between S=0 and 2 states, as no cross terms are found. En, where n = 1, 2, 3, $\dots$, represents the $n$th eigenstate of the Hamiltonian in Eq. (II.2) in the notation used later in this work, incorporating color-spin configuration mixing.
@@
$Y_{c}$ 4400 4431 4446 4560 4580 4615
$Y_{b}$ 10680 10757 10792 10882 10891 10913
Experimental data on $Y_{c}$ and $Y_{b}$ state candidates from different decay processes are listed in Table 5 and will be reviewed and discussed separately below. The assignments of S=0 $Y_{c}$ tetraquarks, along with their theoretical masses and two-body decay branching ratios to $\omega\chi_{c0}$, $\omega\chi_{c1}$, $\omega\chi_{c2}$, and $\rho J/\psi$, are presented in Table 6. Similarly, the assignments of S=0 $Y_{b}$ tetraquarks, along with their theoretical masses and two-body decay branching ratios to $\omega\chi_{b0}$, $\omega\chi_{b1}$, and $\omega\chi_{b2}$, are provided in Table 7.
@@
$\psi(4230)$ Y(4230) $4220$, $46$ Ablikim et al. (2020a) $4225$, $73$ Ablikim et al. (2022a) $4230$, $38$ Ablikim et al. (2015) $4218$, $66.0$ Ablikim et al. (2017) $4219$, $82.7$ Ablikim et al. (2020b)
@@
Y(4380) $4390$, $143$ Ablikim et al. (2021) $4392$, $140$ Ablikim et al. (2017) $4382$, $136$ Ablikim et al. (2020b)
@@
$\Upsilon(10753)$ Y(10753) $10753$, $36$ Mizuk et al. (2019) likely Adachi et al. (2023)
@@
Table 6: Presenting theoretical mass and branching ratio predictions, along with assignments for mixed S=0 $Y_{c}$ tetraquark states. The mass results are given in units of MeV.
En E1 E2 E3 E4 E5 E6 E7 E8 E9 E10 E11 E12 E13 E14 E15 E16
Mass 4209 4258 4338 4363 4369 4435 4466 4480 4509 4522 4534 4542 4592 4617 4650 4673
$\omega\chi_{c0}$ 1 54 4 36 99 9 19 15 77 17 51 7 131 66 4 123
$\omega\chi_{c1}$ 3 161 12 109 297 27 57 45 230 52 154 20 394 198 11 368
$\omega\chi_{c2}$ 5 269 20 181 496 45 94 75 383 86 257 33 657 329 19 613
$\rho J\psi$ 9 484 35 326 892 82 170 136 689 155 462 60 1182 593 34 1103
Data Y(4230) Y’(4230) Y(4340) Y(4380) Y’(4380) Y(4484) Y(4544) Y(4660)
@@
III.2 Y(4230)
In the past two decades, the Y(4230) is an exotic particle observed and confirmed by many experimental collaborations, and is considered part of the broader family of charmonium-like states. First detected in electron-positron annihilation experiments with a mass of approximately 4.23 GeV Aubert et al. (2005), placing it in the region above the open-charm threshold. Unlike conventional quarkonium, which consists of a quark-antiquark pair, the Y(4230) is classified as an exotic hadron due to its unconventional properties, including anomalous decay channels and unexpected production mechanisms. In the previous work, the Y(4230) cannot be accommodated in a conventional charmonium picture including S-D mixing Zhao et al. (2024).
Y(4230) has been observed by many decay processes but mainly from electron positron collision, where the processes include $e^{+}e^{-}\to\pi^{+}\pi^{-}\psi$Ablikim et al. (2020a, 2022b, 2021), $\eta J\psi$Ablikim et al. (2020b), $\omega\chi_{c0}$blikim et al. (2015), $\pi^{+}\pi^{-}h_{c}$blikim et al. (2017), and $K^{+}K^{-}J/\psi$Ablikim et al. (2022a), which are listed in Table 5. There are some clues about states identification appearing when the resonances are observed from different channels.
The resonance parameter of Y(4230) in the invariant mass of $\omega\chi_{c0}$blikim et al. (2015) is inconsistent with the line shape of the Y(4260) observed in $\pi^{+}\pi^{-}J/\psi$Aubert et al. (2005) which is the first observed Y states in 2005 by BABAR. This Y(4260) state later superseded by a updated workLees et al. (2012) in PDG because two interfering resonances are not excluded. But both work were from a single-resonance fit.
Since the Y(4230) was also observed in $K^{+}K^{-}J/\psi$Ablikim et al. (2022a) process, and the resonance parameters is consistent with those from Y(4230) observed from $\pi^{+}\pi^{-}J/\psi$ channel, one may classify the Y(4230) from $K^{+}K^{-}J/\psi$Ablikim et al. (2022a) and $\pi^{+}\pi^{-}J/\psi$ to be a same state.
Later, Y(4230) is observed in the process $\pi^{+}\pi^{-}h_{c}$blikim et al. (2017), whose parameters are different from those of Y(4260) observed in $\pi^{+}\pi^{-}J/\psi$ channel in PDG before year 2015. However, the resonance parameters of Y(4230) in $\pi^{+}\pi^{-}h_{c}$blikim et al. (2017) are consistent with those of the resonance observed in $\omega\chi_{c0}$ Ablikim et al. (2015), one may support that the Y(4230) from $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$ are likely a same state.
The discussion above may support to separate Y(4230) into two different states. In assignment Table 6, the one from $\pi^{+}\pi^{-}J/\psi$ and $K^{+}K^{-}J/\psi$ might be tentatively named Y(4230) while the one observed from $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$ might be tentatively named Y’(4230) respectively. The Y(4230) are assigned to be E1 state of P-wave tetraquark with quantum number $1^{--}$ and Y’(4230) are assigned to be E2 state due to their different decay ratio, where E2 state has much larger ratio for $\omega\chi_{c0}$ channel.
III.3 Y(4360)
As a PDG established state, in PDG data listed in Table 5, the mass of $\psi(4360)$ are from 4.3 to 4.4 GeV, which might show the two different states in the large range. One is around 3.4 GeV, and another is around 3.8GeV, which are renamed as Y(4340) and Y(4380) in the assignment respectively.
The Y(4340) observed only in $\pi^{+}\pi^{-}J/\psi$ process is assigned to be E3 tetraquark state taking a small ratio to $\omega\chi_{c0}$ process which is similar to E1 Y(4230) state, one may be the excitation state in same direction with Y(4230) from $\pi^{+}\pi^{-}J/\psi$ process.
Similar with the discussion for Y(4230), Y(4380) is also observed in $\pi^{+}\pi^{-}J/\psi$ Ablikim et al. (2021) and $\pi^{+}\pi^{-}h_{c}$ Ablikim et al. (2017) respectively, where there line shapes is inconsistent with each other Ablikim et al. (2017). One may support that there are two Y states around the mass 4380MeV, which are tentatively named Y(4380) for the one from $\pi^{+}\pi^{-}J/\psi$, and Y’(4380) for the one from $\pi^{+}\pi^{-}h_{c}$.
In the assignment Table 6, E5 state shows similar property with the Y state from $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$, where E5 state has larger ratio to $\omega\chi_{c0}$ process comparing with E4 state, one may assign Y(4380) and Y’(4380) to be E4 and E5 state respectively.
III.4 Y(4660)
In the assignment Table 6, the Y(4660) is mostly observed in the $\pi^{+}\pi^{-}\psi$ channel, which might be naturally grouped with Y(4230), Y(4340) and Y(4380) from the same process. One may assign it to be E15 tetraquark state which has a small decay ratio to $\omega\chi_{c0}$.
III.5 Other $Y_{c}$ states
[…]
III.6 Y(10753)

### v3

@@ Abstract
The masses of $1^{--}$ P-wave charmonium-like and bottomonium-like tetraquark states are calculated in a constituent quark model (CQM) where the Cornell-like potential and Breit-Fermi interaction are employed. All model parameters were imported from previous work, and predetermined by studying the low-lying conventional S- and P-wave light, charmed, bottom, charmonium, and bottomonium meson mass spectra. The lowest $1^{--}$ tetraquark mass is predicted to be around 4.15 GeV. The decay widths of $1^{--}$ P-wave tetraquark states are calculated for possible two-body strong decay channels within the rearrangement mechanism, including $\omega\chi_{cJ}$, $\eta J/\psi$, and $\rho J/\psi$ for charmonium-like tetraquarks, and $\omega\chi_{bJ}$ for bottomonium-like tetraquarks. The tetraquark theoretical results are compared with the selected exotic states, also known as Y states, and tentative assignments are suggested. This study suggests that $\psi(4230)$, $\psi(4360)$, $\psi(4660)$, and $\Upsilon$(10753) may be P-wave tetraquark states and that multiple states might exist around 4.36 GeV.
@@ II.1 Quark configurations of tetraquark
The construction of tetraquark states follows to the principle that a tetraquark state must be a color singlet, meaning the color wave function of the tetraquark must be a $222]_{1}$ singlet within the $SU_{c}(3)$ group. In this work we consider only charmonium-like and bottomonium-like ($q_{1}Q_{2}\bar{q}_{3}\bar{Q}_{4}$) states, where $q_{1}$ and $Q_{2}$ are light quark and heavy quark, and $\bar{q}_{3}$ and $\bar{Q}_{4}$ are light antiquark and heavy antiquark respectively.
The permutation symmetry of the two-quark cluster ($qQ$) in tetraquark states is described by the Young tabloids $2]_{6}$ and $11]_{\bar{3}}$ of the $SU_{c}(3)$ group, while the color configuration of the two-antiquark cluster ($\bar{q}\bar{Q}$) consists of a $211]_{3}$ triplet and a $22]_{\bar{6}}$ antisextet.
Thus, a $222]_{1}$ color singlet for tetraquark states demands the following configurations: $2]_{6}(q_{1}Q_{2})\otimes[22]_{\bar{6}}(\bar{q}_{3}\bar{Q}_{4})$ and $11]_{\bar{3}}(q_{1}Q_{2})\otimes[211]_{3}(\bar{q}_{3}\bar{Q}_{4})$, which correspond to the color sextet-antisextet ($6_{c}\otimes\bar{6}_{c}$) and triplet-antitriplet ($\bar{3}_{c}\otimes 3_{c}$) configurations, respectively. The explicit color wave function for each tetraquark color configuration can be found in the previous work Zhao et al. (2021).
@@
$\times\psi_{n_{\chi_{1}}l_{\chi_{1}}}(\vec{\chi}_{1}\,)\otimes\psi_{n_{\chi_{2}}l_{\chi_{2}}}(\vec{\chi}_{2}\,)\otimes\psi_{n_{\chi_{3}}l_{\chi_{3}}}(\vec{\chi}_{3}),$
@@ II.2 Hamiltonian
$\vec{L_{ij}}=\vec{r_{ij}}\times\vec{p_{ij}}=\vec{r_{ij}}\times\frac{m_{i}\vec{p_{i}}-m_{j}\vec{p_{j}}}{m_{i}+m_{j}}.$ (8)
@@
By importing the coupling constants $a$, $b$, and $B_{0}$, as well as the constituent quark masses $m_{u}$, $m_{c}$, and $m_{b}$ from previous work Zhao et al. (2021), the mass spectra of S- and P-wave light, charmed, bottom, charmonium, and bottomonium conventional mesons are calculated in the Hamiltonian in Eq. (II.2). A comparison of the theoretical results, presented in Table 1, with experimental data from the PDG Navas et al. (2024), yields a model coupling parameter $\sigma_{0}=0.7$.
@@ II.3 Two-body strong decay
The two-body strong decay properties of the bound states, specifically the Y tetraquark states within the rearrangement mechanism, are studied, as shown in Figure 1.
@@ III.1 Results
The predetermined as well as imported parameters are applied to predict the mass of Y tetraquark states in the Hamiltonian in Eq. (II.2) including the color dependent central potential $V_{cen}$ which may mix up different color-spin configurations.
Because of the cross terms from color operator $(\vec{\lambda}_{i}\cdot\vec{\lambda}_{j})$ and spin operator $(\vec{\sigma}_{i}\cdot\vec{\sigma}_{j})$, listed in Appendix B, eigenstates of the Hamiltonian for S=0 are linear combinations of $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(0\otimes 0)}$, $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=0}_{(1\otimes 1)}$, $\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(0\otimes 0)}$, and $\psi^{c}_{6\otimes\bar{6}}\psi^{S=0}_{(1\otimes 1)}$. Eigenstates for S=2 are linear combinations of $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=2}_{(1\otimes 1)}$ and $\psi^{c}_{6\otimes\bar{6}}\psi^{S=2}_{(1\otimes 1)}$. There is no configuration mixing between S=0 and 2 states, as no cross terms are found. En, where n = 1, 2, 3, $\dots$, represents the $n$th eigenstate of the Hamiltonian in Eq. (II.2) in the notation used later in this work, incorporating color-spin configuration mixing.
@@
$Y_{c}$ 4300 4347 4467 4574 4575 4614
$Y_{b}$ 10568 10665 10791 10828 10851 10941
Experimental data on $Y_{c}$ and $Y_{b}$ state candidates from different decay processes are listed in Table 5 and will be reviewed and discussed separately below. The assignments of S=0 and S=2 $Y_{c}$ tetraquarks, along with their theoretical masses and two-body decay branching ratios to $\omega\chi_{c0}$, $\omega\chi_{c1}$, $\omega\chi_{c2}$, and $\rho J/\psi$, are presented in Table 6. Similarly, the assignments of S=0 $Y_{b}$ tetraquarks, along with their theoretical masses and two-body decay branching ratios to $\omega\chi_{b0}$, $\omega\chi_{b1}$, and $\omega\chi_{b2}$, are provided in Table 7.
@@
$\psi(4230)$ Y(4230) $4220$, $46$ Ablikim et al. (2020a) $4225$, $73$ Ablikim et al. (2022a) $4218$, $28$ Ablikim et al. (2019) $4218$, $66.0$ Ablikim et al. (2017a) $4219$, $82.7$ Ablikim et al. (2020b)
@@
Y(4390) $4390$, $143$ Ablikim et al. (2021) $4392$, $140$ Ablikim et al. (2017a) $4382$, $136$ Ablikim et al. (2020b)
@@
$\Upsilon(10753)$ Y(10753) $10753$, $36$ Mizuk et al. (2019) Yes Adachi et al. (2023)
@@
Table 6: Presenting theoretical mass and branching ratio predictions, along with assignments for mixed S=0 and S=2 $Y_{c}$ tetraquark states. The mass results are given in units of MeV. Y(4390) indicate the states observed in $\pi^{+}\pi^{-}h_{c}$ process. Y’(4390) indicate the states observed in $\pi^{+}\pi^{-}\psi$ process.
Mass 4153 4214 4300 4313 4315 4347 4391
$\omega\chi_{c0}$ 1 54 4 36 99
Data X(4160) Y(4230) Y(4300) Y(4340) Y(4390)
Mass 4413 4467 4488 4496 4518 4536 4562
$\omega\chi_{c0}$ 9 19 15 77 17 51
Data Y’(4390) Y(4484) Y(4484) Y(4544)
Mass 4574 4575 4584 4592 4604 4614 4621
$\omega\chi_{c0}$ 7 131 66 4
Data Y(4660)
III.2 Y(4230)
In the past two decades, the Y(4230) is an exotic particle observed and confirmed by many experimental collaborations, and is considered part of the broader family of charmonium-like states Navas et al. (2024). The Y(4260), the original name of Y(4230), was observed by BaBar in the $\pi^{+}\pi^{-}J/\psi$ invariant-mass spectrum Aubert et al. (2005), and was subsequently confirmed by CLEO Collaboration He et al. (2006) and Belle Collaboration Yuan et al. (2007) in the same process. With improved statistical precision, BESIII identified an asymmetry in the cross section, which caused the peak position to move to a lower mass Ablikim et al. (2017b). The Y(4260) was therefore renamed Y(4230).
Unlike conventional quarkonium, which consists of a quark-antiquark pair, the Y(4230) is classified as an exotic hadron due to its unconventional properties, including anomalous decay channels and unexpected production mechanisms. In the previous work, the Y(4230) cannot be accommodated in a conventional charmonium picture including S-D mixing Zhao et al. (2024).
[…]
The Y(4230) was also observed in $K^{+}K^{-}J/\psi$ Ablikim et al. (2022a) process, and the resonance parameters were consistent with those from Y(4230) observed from $\pi^{+}\pi^{-}J/\psi$ channel, one may classify the Y(4230) from $K^{+}K^{-}J/\psi$ and $\pi^{+}\pi^{-}J/\psi$ to be a same state. Later, the Y(4230) was observed in the process $\pi^{+}\pi^{-}h_{c}$ Ablikim et al. (2017a), whose resonance parameters were consistent with those of the resonance observed in $\omega\chi_{c0}$ Ablikim et al. (2015), one may support that the Y(4230) from $\omega\chi_{c0}$ and $\pi^{+}\pi^{-}h_{c}$ are likely a same state.
The number of states in this mass region remains to be further determined experimentally Navas et al. (2024). This study predicts only one state in the mass region. As shown in Table 6, we may assign the Y(4230), observed in the $\pi^{+}\pi^{-}J/\psi$, $K^{+}K^{-}J/\psi$, $\omega\chi_{c0}$, and $\pi^{+}\pi^{-}h_{c}$ channels, to be the second tetraquark state, which is based on its large decay ratio to $\omega\chi_{c0}$, with a corresponding theoretical mass of 4214 MeV.
III.3 Y(4360)
As a PDG established state, in PDG data listed in Table 5, the mass of $\psi(4360)$ are from 4.3 to 4.4 GeV, which might show the three different states in the large range. The first state is around 4.30 GeV, the second state is around 4.34 GeV, and the third state is around 4.39 GeV, which are renamed as Y(4300), Y(4340), and Y(4390) in the assignment respectively.
The Y(4300) and Y(4340) observed in $\pi^{+}\pi^{-}J/\psi$ and $\pi^{+}\pi^{-}\psi(2S)$ process are assigned to be S=2 tetraquark state, with a corresponding theoretical mass of 4300 MeV and 4347 MeV respectively.
The Y(4390) was observed in $\pi^{+}\pi^{-}J/\psi$ Ablikim et al. (2021) and $\pi^{+}\pi^{-}h_{c}$ Ablikim et al. (2017a) channels respectively. In the $\pi^{+}\pi^{-}h_{c}$ process Ablikim et al. (2017a), the Y(4230) was also observed. The resonance parameters of the Y(4230) in $\pi^{+}\pi^{-}h_{c}$ process were consistent with those of the Y(4230) observed in $\omega\chi_{c0}$ Ablikim et al. (2015), which may suggest that the Y(4390) could also be observed in the $\omega\chi_{c0}$ channel.
In the assignment Table 6, the Y(4390) observed in the $\pi^{+}\pi^{-}h_{c}$ channel is assigned based on its large decay ratio to $\omega\chi_{c0}$, with a corresponding theoretical mass of 4391 MeV. The Y(4390) could potentially be observed in the $\omega\chi_{c0}$ channel, which may motivate future experimental searches.
In Ref Ablikim et al. (2017a), the line shapes of Y(4390) observed in $\pi^{+}\pi^{-}h_{c}$ process Ablikim et al. (2017a) are inconsistent with those in $\pi^{+}\pi^{-}J/\psi$ and $\pi^{+}\pi^{-}\psi(2S)$ by BaBar Aubert et al. (2005, 2007); Lees et al. (2014, 2012) and by Belle Yuan et al. (2007); Wang et al. (2007, 2015); Liu et al. (2013). One may support that there are two Y states around the mass 4390MeV, which are tentatively named Y(4390) for the one from $\pi^{+}\pi^{-}h_{c}$, and Y’(4390) for the one from $\pi^{+}\pi^{-}J/\psi$.
In the assignment Table 6, the Y(4390) observed in the $\pi^{+}\pi^{-}\psi$ channel is assigned based on the small decay ratio to $\omega\chi_{c0}$, with a corresponding theoretical mass of 4413 MeV.
III.4 Y(4660)
In the assignment Table 6, the Y(4660) is mostly observed in the $\pi^{+}\pi^{-}\psi$ channel, which might be naturally grouped with Y(4230), Y(4300), Y(4340) and Y’(4390) from the same process. One may assign it to be tetraquark state which has a small decay ratio to $\omega\chi_{cJ}$, with a corresponding theoretical mass of 4621 MeV.
III.5 Other $Y_{c}$ states
The $X(4160)$ was first observed by the Belle Collaboration in double-charmonium production, $e^{+}e^{-}\to J/\psi X(4160)$ with $X(4160)\to D^{+}D^{-}$Pakhlov et al. (2008). Later, a related state was observed by the LHCb Collaboration in the decay $B^{+}\to J/\psi\phi K^{+}$Aaij et al. (2021), reconstructed from the invariant mass spectrum of $J/\psi\phi$. The two states observed around 4.15 GeV are grouped together as a single entry in the PDG Navas et al. (2024), according to their similar masses and decay widths. Whether these observations correspond to the same state still requires further experimental confirmation. The nature of the $X(4160)$ therefore remains uncertain and requires further theoretical and experimental investigation.
The production rate of the $X(4160)$ in double-charmonium processes, along with its decay patterns into the $D^{+}D^{-}$ and $J/\psi\phi$ channels, does not align well with expectations for conventional $c\bar{c}$ states Chao (2008); Navas et al. (2024). But, some studies have still considered interpretations of the $X(4160)$ as a higher charmonium state Yang et al. (2010). The $X(4160)$ is more likely interpreted as an exotic hadron, with several theoretical scenarios proposed. These include a compact tetraquark configuration, either hidden-charm $c\bar{c}q\bar{q}$ or $c\bar{c}s\bar{s}$, supported by QCD sum rules Chen et al. (2017), as well as a meson-meson molecular state, in particular a $D_{s}^{*}\bar{D}_{s}^{*}$ bound state Wang et al. (2018).
Despite these scenarios, the quantum numbers of the $X(4160)$ remain unclear. The Belle Collaboration did not measure its quantum numbers Pakhlov et al. (2008), while LHCb analyses favor an assignment of $J^{PC}=2^{-+}$ Aaij et al. (2021). Given its measured mass near 4.16 GeV, the $X(4160)$ may also be interpreted as a tetraquark candidate, based on mass consistency with theoretical predictions for hidden-charm exotics in this region Chen et al. (2017). However, further experimental confirmation is still required. In the assignment, $X(4160)$ is tentatively assigned to be tetraquark state due to masses agreement, with a corresponding theoretical mass of 4153 MeV.
[…]
III.6 Y(10753)